OpenStudy (anonymous):

4/5 (25/12)....i was also wondering if anyone has good ideas on studying for math

OpenStudy (anonymous):

Just multiply straight across the numerator and the denominator. \[{4 \over 5} \cdot {25 \over 12}\]\[{4 \over 5} {\rightarrow \over \rightarrow} {25 \over 12} = {4 \cdot 25 \over 5 \cdot 12}\] Do you need to write in lowest terms? Math is different to study. It's move than just memorization. I would suggest practice, practice, practice. And try to think about things in a way that makes sense to you. Pneumonic's help. Of course, OpenStudy is a great study resource! :-)

OpenStudy (anonymous):

Thank you! and thank you for the tips on studying